OpenStudy (anonymous):

Solve: -2x - 15 = 6x + 9 x= -3 right?

OpenStudy (austinl):

\(-2x - 15 = 6x + 9\) \(-2x-6x=9+15\) \(-8x=24\) \(x=-3\) Correct.

OpenStudy (anonymous):

yes :)